A guitar string has a linear density of 8.30*10^-4 kg/m and a length of 0.660m. The tension in the string is 52.0N. When the fundamental frequency of the string is sounded with a 196.0Hz tuning fork, what beat frequency is heard?
Answer #1

L=0.660m
Ft=52.0 N
myuu=8.3*10^-4 kg/m

f1=1/2L(Ft/myuu)^1/2
f1=1/2(0.660m)(52/8.3*10^-4)
f1=189.62 Hz

To find beat frequency:
fbeat=f1-f2
fbeat=189.62 Hz -196.0 Hz= -6.38 Hz

=6.38 HZ ANSWER